# Toward Self-Improvement of LLMs via Imagination, Searching, and Criticizing

### v1

@@ 4.4 Critic
ORM
In additional to the value function and PRM, we introduce ORM to guide MCTS. ORM is designed to evaluate options sequences in their entirety, assessing the extent to which the complete trajectory aligns with the desired end goal. The outcome evaluation complements value function and PRM by offering a comprehensive assessment of trajectories. Crucially, ORM plays a vital role in the simulation stage of MCTS by providing more accurate signals on the terminal state, which in turn facilitates a more balance between exploration and exploitation strategies. ORM is formulated as a text generation task, similar to PRM. We leverage the same dataset for the value function training and construct ${\mathcal{D}}_{\mathtt{ORM}}=\{({\bm{x}}_{i},{\bm{o}}_{1:T}^{i},r_{i}^{\mathtt{ORM}})|i\in[N]\}$, where each instance includes a initial state or prompt ${\bm{x}}_{i}$, a sequence of actions or options ${\bm{o}}_{1:T}^{i}$ taken from that state, and a textual reward $r_{i}^{\mathtt{ORM}}$ indicating the sequence’s success or quality. Similarly, ORM is initialized from the policy model $\pi$ and the following prompt templates and language model loss are used for training.
[…]
4.5 Policy Self-Improvement
We have discussed how $\eta$Mcts can guide policy to find trajectories of higher quality and. In this subsection, we discuss how to leverage these trajectories to further improve the policy. It is an iterative process with each iteration containing two main steps: data generation and policy finetuning.
Data generation
In this step, we assume to have the current policy $\pi_{\theta_{k}}$ and synthetic prompts ${\mathcal{D}}_{k}=\{{\bm{x}}^{k}_{1},\dots\}$ at the $k$-th round, where each ${\bm{x}}^{k}_{1}$ represents a question. We obtain the corresponding training data ${\mathcal{D}}_{k}$ for policy $\pi_{\theta_{k}}$ by firstly performing $\eta$Mcts on ${\mathcal{D}}_{k}$4.3) and then sampling a trajectory ${\bm{y}}^{k}_{i}$ from the corresponding MCTS forest for each question ${\bm{x}}^{k}_{i}$. There are several ways to select a trajectory from a MCTS forest, such as taking a greedy path based on the critic score ($w_{i}$ in Eq. 1). Here we choose the trajectory that yield the highest critic score on the leaf node for each input question. As the next step, we filter out instances where the corresponding trajectory is not in high quality:
${\mathcal{D}}_{k}=\{({\bm{x}}^{k}_{i},{\bm{y}}^{k}_{i})~{}|~{}f({\bm{x}}^{k}_{i},{\bm{y}}^{k}_{i})>\gamma\}$
where $f$ represents the quality evaluating function for quality scoring, and $\gamma$ represents the threshold. There can be several ways to implement the function, and here we simply use the ORM4.4).
Policy finetuning
With the obtained training data ${\mathcal{D}}_{k}$, we organize the data into the following prompt templates:
[…]
Then the policy $\pi_{\theta_{k}}$ is finetuned using target-loss SFT:
$\mathcal{L}_{\theta_{k}}=\mathbb{E}_{({\bm{x}}^{k}_{i},{\bm{y}}^{k}_{i})\sim{\mathcal{D}}_{k}}\big{[}\log\pi_{\theta_{k}}({\bm{y}}^{k}_{i}|{\bm{x}}^{k}_{i})\big{]}$
This results in an updated policy $\pi_{\theta_{k+1}}$. We leave other training methods, such as DPO (Rafailov et al., 2023) or PPO (Schulman et al., 2017) in future work.
5 Experiments
5.1 Evaluation Setups
Datasets
AlphaLLM is generally applicable to a wide spectrum tasks. As an early exploration, in this paper, we conduct experiments on mathematical reasoning problems where the learning signals are clear to define i.e., , final answer is correct or wrong. We choose to evaluate on two widely used datasets GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). For GSM8K, we utilize the whole test set while for MATH, due to computation constraints, we utilize a subset following the same procedure of Lightman et al. (2023).
Metrics
We evaluate the performance of predicting answers correctly for policy models. In the same time, we calculate the average rollouts, represented by the number of nodes in the tree, as a measure of computational efficiency.
5.2 Baseline Systems
We evaluate the performance of AlphaLLM against a suite of proprietary model, including OpenAI’s GPT-4 and GPT-3.5, Anthropic’s Claude-2, as well as Google’s PaLM-2 and the gemini model family. To ensure a fair and consistent evaluation, we employ CoT as our primary prompting method. We additionally report PAL (Gao et al., 2023) prompting performance with GPT-4 as it demonstrates enhanced performance.
Additionally, we conduct comparisons with strong open-source models, including LLaMA-2 70B (Touvron et al., 2023a) and Wizardmath 70B (Luo et al., 2023). For LLaMA-2 70B, we present results from few-shot prompting as well as zero-shot prompting for its SFT version, which was trained using CoT rationales and final answers. Wizardmath 70B has been trained on a diverse set of mathematical data generated by ChatGPT, employing both SFT and RLHF. We provide zero-shot prompting results.
5.3 Implementation Details
We select LLaMA-2 70B as the policy model for the GSM8K dataset and Wizardmath 70B V10 for the MATH dataset. To construct the training dataset for the value function, PRM and ORM, we generate 50 trajectories for each prompt and construct the training target following Section 4.4. Both PRM and ORM are initialized using the weights from the policy model. In the design of ORM, tool usage is not incorporated for GSM8K. However, for MATH, we enhance ORM by incorporating tools like pythoin sympy to assess the quality of a trajectory, in a manner similar to that described by Gou et al. (2023b). The training employ a learning rate of 1e-6 and are trained for one epoch. For the fast rollout policy model, we opt for the Abel-002-7B model (Chern et al., 2023) for both the GSM8K and MATH tasks for its high efficiency and superior performance.
We set the MCTS parameters as follows: in GSM8K, $c=1$ for the small scale (#rollout) and $1.5$ for the large scale, with $\alpha=1$. For $t=0$, $c_{\text{min}}(0)=10$ for the small scale and $40$ for the large scale, while for the rest of $t$, $c_{\text{min}}(t)=2$. We also set $c_{\text{max}}(0)=10$ for the small scale and $40$ for the large scale, and for the remaining $t$, $c_{\text{max}}(t)=10$. The termination condition is based on sentence termination. In MATH, the parameters are $c=1$, $\alpha=1$, and for $t=0$, $c_{\text{min}}(0)=10$ for the small scale and $20$ for the large scale, while for the rest of $t$, $c_{\text{min}}(t)=3$. We set $c_{\text{max}}(0)=10$ for the small scale and $20$ for the large scale, and for the remaining $t$, $c_{\text{max}}(t)=10$. The termination function is rule-based, checking if there are any formulations or calculations in the sentence. If there are, the option is terminated; otherwise, the option continues to extend.
For policy self-improving (§4.5), we train the policy model up to 3 epochs, setting batch size to 128, learning rate to $5\times 10^{-6}$ and minimal learning rate to $1\times 10^{-6}$. Linear warm-up and decay is used with warm-up percent to be 10%. We perform early stopping based on a devset held out from the training instances. For second-round self-improving, we sample 7.9k MetaMath (Yu et al., 2023) prompts to obtain the corresponding MCTS outputs for training.
@@ 5.4 Results
LLaMA-2 70B Greedy 0 $\times$ $\times$ $\times$ 57.8 -
LLaMA-2 70B SFT Greedy 7.5k $\checkmark$ $\checkmark$ $\times$ 69.3 -
WizardMath 70B V1.0 Greedy 96k $\checkmark$ $\checkmark$ $\times$ - 20.7
AlphaLLM Greedy 7.5k/3k $\times$ $\checkmark$ $\checkmark$ 73.7 23.6
AlphaLLM $\eta$Mcts 7.5k/3k $\times$ $\checkmark$ $\times$ 88.9 48.7
AlphaLLM $\eta$Mcts 7.5k/3k $\times$ $\checkmark$ $\checkmark$ 92.0 51.0
Table 2: Comparison results of AlphaLLM on the GSM8K and MATH datasets, utilizing LLaMA-2 70B and WizardMath 70B V1.0 as base models for GSM8K and MATH datasets, respectively. #Annotation indicates the quantity of labeled data employed for fine-tuning each base model. The annotation used for training are noted as RN for rationales and FA for final answers. SYN means models trained on synthetic prompts, where trajectories were generated using $\eta$Mcts.
Table 2 lists the performance comparisons of various methods on the GSM8K and MATH datasets. Our findings reveal that AlphaLLM, which utilizes only final answer annotations and self-improves through the training on synthetic prompts with responses from $\eta$Mcts, outperforms both LLaMA-2 70B and WizardMath 70B V1.0—even though these models are trained on a larger set of examples that include both rationales and final answer annotations. This comparison underscores the efficacy and broad applicability of our imagination-searching-criticizing self-improving framework. Moreover, when our model is augmented with $\eta$Mcts decoding strategy, its performance markedly improves, achieving scores of 88.9 and 48.7 on the GSM8K and MATH datasets, respectively. Following two iterations of self-improvement using synthetic prompts, AlphaLLM demonstrates performance comparable to that of GPT-4. This suggests a viable approach to improving LLMs’ capabilities in complex problem-solving tasks in a self-improving fashion, leveraging a minimal amount of labeled data.
In addition, table 3 presents the performance of various methods applied to different number of responses, from 10 to 50. Our analysis confirms several key findings: 1) Reranking utilizing ORM consistently outperforms self-consistency techniques, indicating that ORM is capable of generating meaningful signals for searching. 2) $\eta$Mcts demonstrates superior performance while requiring significantly fewer rollouts. For instance, on the MATH dataset, $\eta$Mcts achieves better results with only half the number of rollouts compared to reranking. These results suggest that our design of an efficient MCTS in AlphaLLM can serve as an effective policy improvement operation, enabling the search for high-quality trajectories with reduced computational cost.

### v2

@@ 5.2 Results
Llama-2-70b Greedy 0 $\times$ $\times$ $\times$ 57.8 -
Llama-2-70b SFT Greedy 7.5k $\checkmark$ $\checkmark$ $\times$ 69.3 -
WizardMath-70B-V1.0 Greedy 96k $\checkmark$ $\checkmark$ $\times$ - 20.7
AlphaLLM Greedy 7.5k/7.5k $\times$ $\checkmark$ $\checkmark$ 73.7 23.6
AlphaLLM $\eta$Mcts 7.5k/7.5k $\times$ $\checkmark$ $\times$ 88.9 48.7
AlphaLLM $\eta$Mcts 7.5k/7.5k $\times$ $\checkmark$ $\checkmark$ 92.0 51.0
Table 2: Comparison results of AlphaLLM on the GSM8K and MATH datasets. #Annotation indicates the quantity of labeled data employed for fine-tuning policy or training critic models. The annotation used for training are noted as RN for rationales and FA for final answers. SYN means models trained on synthetic prompts, where trajectories were generated using $\eta$Mcts.
Table 2 lists the performance comparisons of various methods on the GSM8K and MATH datasets. Our findings reveal that AlphaLLM, based on Llama-2-70B and WizardMath-70B-V1.0, utilizes only final answer annotations and continues to improve through training on responses from $\eta$Mcts. This comparison underscores the efficacy and broad applicability of our imagination-searching-criticizing self-improving framework. Moreover, when our model is augmented with $\eta$Mcts decoding strategy, its performance markedly improves, achieving scores of 88.9 and 48.7 on the GSM8K and MATH datasets, respectively. Following two iterations of self-improvement using synthetic prompts, AlphaLLM demonstrates performance comparable to that of GPT-4. This suggests a viable approach to improving LLMs’ capabilities in complex problem-solving tasks in a self-improving fashion, leveraging a minimal amount of labeled data. We also analyze the performance of various search methods in Appendix A.8.
@@ 5.3 Ablation Study
AB PRM FR-ORM SM LG-#Rollout Acc
$\times$ $\times$ $\times$ $\times$ $\times$ 79.5
$\checkmark$ $\times$ $\times$ $\times$ $\times$ 84.9
$\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ 85.9
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ 86.5
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 87.0
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 88.9
(a) Ablation study on GSM8K
TA-ORM Option Acc #Rollout
$\times$ $\times$ 38.8 201
$\checkmark$ $\times$ 44.1 198
$\checkmark$ $\checkmark$ 45.4 148
(b) Ablation study on MATH
Table 3: (a): Ablation studies on the GSM8K test set of various components of $\eta$Mcts, including adaptive branching, PRM, fast-rollout with ORM, state merge, and large number of rollouts. (b): Ablation studies of the impacts of tool-augmented ORM and option-level formulation on MATH.
We assess the effectiveness of each component in AlphaLLM and report the results on GSM8K in Table 3(a). Vanilla MCTS, configured with only the value function and a fixed number of children per node, achieves an accuracy of 79.5%. This serves as a reference point for evaluating the incremental benefits introduced by each additional component. The use of adaptive branching increae the accuracy to 84.9%. The addition of PRM improves the accuracy modestly to 85.9%, showing the effectivenss of process supervision for searching. A more significant improvement is observed with the introduction of ORM with fast rollout, which boosts the accuracy to 86.5%. Integrating state merging results in a further increase in accuracy, reaching 87.0%. Finally the combined of increasing the number of rollouts with the other components yields the best performance on this task.
Figure 2: Empirical analysis on GSM8K of different self-improving data collection methods and number of iterations. Models are evaluated with greedy decoding, $\eta$Mcts with small #rollout and large #rollout.
Figure 2 depicts a comparative results on GSM8K of two rounds of self-improving trained on trajectories collected using reranking and $\eta$Mcts. We report the performance of greedy decoding, $\eta$Mcts with a relatively small number of rollouts (50-60), and $\eta$Mcts with a larger number of rollouts (200-300) for each model. We observe that 1) Models trained on the trajectories from reranking or $\eta$Mcts outperform the initial policy by a significant margin. In addition, the performance can be iteratively improved with training suggesting that self-improving has the potential to achieve continual performance gain. 2) While both reranking and $\eta$Mcts can generate high-quality trajectories for self-improving , $\eta$Mcts is performant with high efficiency and better accuracy. Models trained on trajectories generated by it not only exceed the performance of those trained on reranked trajectories but also, when decoded with $\eta$Mcts, demonstrate on par performance with GPT-4, revealing that AlphaLLM is an effective self-improving framework.
Method Threshold Acc
Edit distance $20$ $86.8$
Edit distance $50$ $87.0$
Cosine Similarity $0.7$ $86.3$
Model-based N/A $86.7$
(a) Ablation on the choice of state merge functions.
#Trajetory Acc
$1$ $85.9$
$4$ $86.5$
$8$ $86.7$
(b) Ablation on the number of trajectories.
Table 4: (a): Ablation studies on the choice of heuristic/model-based functions in state merge on GSM8K with base Llama2-70b. The model used in the model-based state merge is Llama-2-70b-chat. (b): Ablation studies of the number of rollout trajectories in fast-rollout estimation on GSM8K with base Llama2-70b.
We further analyze the impact of different hyperparameters and design choices for each component. Table 4(a) shows that varying heuristic functions (with hyperparameters) for state merge has limited impact on performance. Table 4(b) shows that, as the number of fast-rollouts increases, there is a corresponding improvement in performance. This is due to the reduction in the variance of the estimates. We used $n=4$ in our experiments for better trade-off between performance and efficiency. Additional ablations on the choice of fast-rollout models, are provided in Appendix A.7.
6 Conclusion
Appendix A Appendix
A.1 Imagination, Searching, Criticizing and Learning Loop
Output $\theta^{k}$
for $k\leftarrow 1,\dots,K$ do
      Construct dataset ${\mathcal{D}}^{k}=\{({\bm{x}}^{k},\hat{{\bm{y}}}^{k})\}$
end for
Algorithm 1 LLM self-improving loop
The algorithm is shown in Algorithm 1.
A.2 Option-level MCTS
As illustrated in Figure 3, option-level MCTS consists of the following operations:
  • Selection Starting from the root node, we iteratively select the child node based on Equation LABEL:eqs:ucb.
  • Simulation The scaled reward of the newly expanded node, as well as some simulated future trajectories are evaluated using the feedback functions, which is discussed in §4.4.
A.3 Importance-Based Adaptive Branching Under Uniform Distribution
Let $V=\{v_{\phi}^{\pi}({\bm{s}}_{t},{\bm{o}}_{t}^{1}),v_{\phi}^{\pi}({\bm{s}}_{t},{\bm{o}}_{t}^{2}),...,v_{\phi}^{\pi}({\bm{s}}_{t},{\bm{o}}_{t}^{m_{t}})\}$ be a set of $m_{t}$ values that are uniformly distributed. If the maximum and minimum values from $V$ are $v_{\max}$ and $v_{\min}$, the average gap between two consecutive values is given by $\frac{v_{\max}-v_{\min}}{m_{t}-1}$. The upper bound of expected minimum distances from a new value $v_{\text{new}}$ to any value from $V$ is achieved when $v_{\text{new}}$ is consistently positioned at the midpoint between two consecutive values, and it is given by $\frac{v_{\max}-v_{\min}}{2(m_{t}-1)}$.
Since $v_{\max}-v_{\min}=2I({\bm{s}}_{t})$ for a uniform distribution, we can conclude that $E_{\phi}(t)\leq\frac{I({\bm{s}}_{t})}{m_{t}-1}$.
Theorem 4.1.
The optimal branching factor $m_{t}$ in a tree search is set such that $m_{t}-1$ is proportional to the node importance $I({\bm{s}}_{t})$, under the condition $\frac{I({\bm{s}}_{t})}{m_{t}-1}\leq\epsilon$.
Proof.
We can have the optimization problem as:
minimize: $\sum m_{t}$
subject to: $\frac{I({\bm{s}}_{t})}{m_{t}-1}\leq\epsilon$
Introduce the Lagrange multiplier $\lambda_{t}$ for each constraint:
$L(m_{t},\lambda_{t})=\sum m_{t}+\sum\lambda_{t}\left(\epsilon(m_{t}-1)-I({\bm{s}}_{t})\right)$
Now, let’s find the gradient of the Lagrangian with respect to $m_{t}$ and $\lambda_{t}$ and set them to zero:
$\nabla_{m_{t}}L$ $=1+\epsilon\lambda_{t}=0$
$\nabla_{\lambda_{t}}L$ $=\epsilon(m_{t}-1)-I({\bm{s}}_{t})=0$
From the first equation, we get:
$\lambda_{t}=-\frac{1}{\epsilon}$
Substitute this value of $\lambda_{t}$ into the second equation:
$\epsilon(m_{t}-1)-I({\bm{s}}_{t})=0$
Solving for $m_{t}$, we get:
$m_{t}=\frac{I({\bm{s}}_{t})}{\epsilon}+1$
Thus, $m_{t}-1$ is proportional to the node importance $I({\bm{s}}_{t})$. ∎
A.4 Importance-Based Adaptive Branching Under Gaussian Distribution
If we assume that $v_{\phi}^{\pi}([{\bm{s}}_{t},{\bm{o}}_{t}^{j}])$ and $v_{\phi}^{\pi}([{\bm{s}}_{t},{\bm{o}}_{t}^{i}])$ are independent and identically distributed Gaussian random variables:
$v_{\phi}^{\pi}([{\bm{s}}_{t},{\bm{o}}_{t}^{j}]),v_{\phi}^{\pi}([{\bm{s}}_{t},{\bm{o}}_{t}^{i}])\sim\mathcal{N}(\mu,\sigma^{2})$
The difference $D_{ij}=v_{\phi}^{\pi}([{\bm{s}}_{t},{\bm{o}}_{t}^{j}])-v_{\phi}^{\pi}([{\bm{s}}_{t},{\bm{o}}_{t}^{i}])$ will follow a normal distribution with:
$D_{ij}\sim\mathcal{N}(0,2\sigma^{2})$
To find the expected minimum absolute difference between $v_{\phi}^{\pi}([{\bm{s}}_{t},{\bm{o}}_{t}^{j}])$ and the closest $v_{\phi}^{\pi}([{\bm{s}}_{t},{\bm{o}}_{t}^{i}])$, we need to consider the distribution of the minimum of $m_{t}$ Gaussian differences.
The expected minimum value of $m_{t}$ absolute differences can be approximated using properties of order statistics for Gaussian distributions.
For a set of $m_{t}$ independent normal random variables with variance $2\sigma^{2}$, the expected minimum absolute difference, $\mathbb{E}[\min_{i}|D_{ij}|]$, can be approximated by:
$E_{\phi}(t)\approx\frac{\sigma\sqrt{2}}{\sqrt{m_{t}}}$
This approximation arises from the fact that the expected minimum value of the absolute deviations of normally distributed random variables scales with the inverse of the square root of the number of samples.
Then, assume the range of the $m_{t}$ samples are $R_{m}=max(v_{\phi}^{\pi}([{\bm{s}}_{t},{\bm{o}}_{t}^{i}])-min(v_{\phi}^{\pi}([{\bm{s}}_{t},{\bm{o}}_{t}^{i}])$, the the expected range $\mathbb{E}[R_{m}]$ of $m_{t}$ samples from a normal distribution can be approximated using properties of extreme values of Gaussian distributions. The range $R_{m}$ can be approximated as:
$R_{m}\approx\sigma(z_{0.9995}-z_{0.0005})$
where $z_{p}$ is the p-th percentile of the standard normal distribution. It can converge to
$R_{m}\approx\sigma\sqrt{2\ln(m_{t})}\left(2-\frac{\ln(\ln(m_{t}))}{4\ln(m_{t})}\right)$
For simplicity, we can approximate the range using the primary term, which captures the dominant behavior:
$R_{m}\approx\sigma\sqrt{2\ln(m_{t})}$
Then we have
$E_{\phi}(t)\approx\frac{\sqrt{2}}{{\sqrt{m_{t}}}}\frac{R_{m}}{\sqrt{2\ln(m_{t})}}$
Knowing that for all distributions,
$I({\bm{s}}_{t})\geq\frac{R_{m}}{2}$
We have
$E_{\phi}(t)\leq\frac{I(s_{t})}{\sqrt{m_{t}\ln(m_{t})}}$
Then to find the optimal $m_{t}$, the optimization problem is
minimize: $\sum m_{t}$
subject to: $\frac{I(s_{t})}{\sqrt{m_{t}\ln(m_{t})}}\leq\epsilon$
To solve this optimization problem, we can first rewrite the constraint in terms of $m_{t}$.
$m_{t}\ln(m_{t})\geq\frac{I^{2}(s_{t})}{\epsilon^{2}}$
Now, let’s define a new function $g(m_{t})=m_{t}\ln(m_{t})$. We want to find the minimum $m_{t}$ such that $g(m_{t})\geq\frac{I^{2}(s_{t})}{\epsilon^{2}}$. To do this, we can find the derivative of $g(m_{t})$ and set it to zero to find the critical points.
$g^{\prime}(m_{t})=\frac{d}{dm_{t}}(m_{t}\ln(m_{t}))=\ln(m_{t})+1$
Setting the derivative to zero:
$\ln(m_{t})=-1$
$m_{t}=e^{-1}$
However, this critical point corresponds to a minimum of the function $g(m_{t})$, and we are interested in the minimum $m_{t}$ that satisfies the constraint $g(m_{t})\geq\frac{I^{2}(s_{t})}{\epsilon^{2}}$. Since the function $g(m_{t})$ is increasing for $m_{t}>e^{-1}$, we can find the minimum $m_{t}$ by setting $g(m_{t})=\frac{I^{2}(s_{t})}{\epsilon^{2}}$ and solving for $m_{t}$:
$m_{t}\ln(m_{t})=\frac{I^{2}(s_{t})}{\epsilon^{2}}$
This can not be solved directly, but we can still observe that there is a positive correlation between $m_{t}$ and $I({\bm{s}}_{t})$.
A.5 Prompt Templates
A.5.1 PRM
[…]
A.5.2 ORM
[…]
A.5.3 Policy Finetuning
For MATH experiments that take a WizardMath V1.0 70B as the policy, we adopt their proposed system prompt for self-improving. For GSM8K experiments taking Llama2 70B pretrain as the policy, we use the following system prompt.
[…]
A.6 MCTS Details
We set the MCTS parameters in Table 5.
Method GSM8K MATH
Small Large Small Large
$c$ 1.0 1.5 1.0 1.0
$\alpha$ 1.0 1.0 1.0 1.0
$c_{\text{max}}(0)$ 60 60 60 60
$c_{\text{max}}(t)$ where $t>0$ 10 10 10 10
$c_{\text{min}}(0)$ 10 40 10 20
$c_{\text{min}}(t)$ where $t>0$ 2 2 3 3
Table 5: Parameters for MCTS. The Small/Large means small #rollout and small #rollout
A.7 Additional Ablations
Fast-rollout model
Using Llama-2-70b instead of Abel-7B-002 improves performance by reducing bias from a smaller model, but Abel-002-7B is faster with similar computational resources due to higher concurrency and quicker processing. The details can be found in Table 6.
Model Acc (%) Speed (s)
Abel-002-7B 87.0 16.8
Llama-2-70B 87.3 38.1
Table 6: Ablation study over different fast-rollout models on GSM8K.
A.8 Search Comparison